OpenStudy (jokersmart):

Check my answer? A crater shaped like a cone has a height of 24 yd and a diameter of 72 yd. What is the volume of the crater? Use 3.14 to approximate pi and express your final answer in hundredths. ___yd3 I think 32555.52

OpenStudy (jdoe0001):

\(\large { \textit{volume of a cone}=\cfrac{\pi r^2 h}{3}\qquad \begin{cases} r\to radius\to \frac{diameter}{2}\\ h\to height \end{cases} }\)